OpenStudy (anonymous):

what does x+3 over 4 - x-4 over 2 = 3 over 8 i got -19 over 2

OpenStudy (campbell_st):

well but each fraction over a denominator of 8 \[\frac{2(x + 3)}{8} - \frac{4(x - 4)}{8} = \frac{3}{8}\] so you are really solving 2(x + 3) - 4( x - 4) = 3 distribute, collect like terms and then solve for x

OpenStudy (anonymous):

so it is a positive 19 over 2?

OpenStudy (campbell_st):

so its - 2x + 22 = 3 - 2x = -19 x = 19/2 yep... thats the answer... well done